a 50

b 4

c 36

d need more info

A population has a mean of 50 and a standard deviation of 12. For a sample size of 4, what is the standard error?

a.12

b.4

c.3

d.6

A population has a mean of 100 and a standard deviation of 15. For a sample size of 25, what is the standard error for the distribution of sample means?

a.2

b.15

c.25

d.3
Answer #1

1.

Mean of sample remains same as the population mean.

So expected value of sample mean = 50

.correct option - (a)

2.

Standard error = std. Deviation / sqrt(sample size)

= 12/sqrt(4) = 12/2 =6

Correct option--(d)

3.

Standard error = 15/sqrt(25) = 15/5= 3

Correct option ---(d)